Circular No. 5819

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

YY HERCULIS U. Munari and L. M. Buson, Asiago and Padova Observatories; R.

J. Ivison, University of Toronto; and G. Sostero and A. Lepardo, Associazione Friulana di Astronomia e Meteorologia (AFAM), Udine, Italy, communicate: “Near-simultaneous ultraviolet, optical, and near-infrared observations have been performed on the symbiotic star YY Her, which is currently experiencing a bright phase (cf. IAUC 5805). Photoelectric photometry with the AFAM 0.45-m reflector on June 15.94 and 18.03 UT gives, respectively, V = 11.67, B–V = +0.30 and V = 11.68, B–V = +0.20, U–B = –0.58 (± 0.04 in all colors), compared to V ∼ 14.2 and B–V ∼ +1.5 in the fall of 1992. Flux calibrated CCD spectra (ranges 400-620 nm with the échelle spectrograph and 350-1050 nm with the Boller & Chivens spectrograph) were secured on June 6 and 10 at the Asiago Observatory’s 1.82-m telescope. Additional spectra were obtained on June 11 with the Manchester échelle spectrograph on the 2.5-m Isaac Newton Telescope in the Canary Islands. Finally, International Ultraviolet Explorer SWP and LWP low-resolution spectra were collected on June 14 at the VILSPA station of the European Space Agency. The optical region, dominated by TiO bands in 1992, is now dominated by hot continuum emission (extending well into the red) and also by an intense Balmer continuum in emission. Compared to the fall of 1992, the absolute intensity of H I, He I, and He II show an ∼ 10-fold increase, while the Hβ/Hγ ratio stayed fixed. [O III] and [N III] are now detected in emission. Preliminary inspection of the Asiago high-resolution spectra does not show signs of relevant mass loss or significant changes in the line profiles. The IUE spectra show 5-fold and 4-fold increases in the continuum at 145.0 and 250.0 nm compared with 1980 May archival data. The He II 164.0-nm line is now by far the strongest emission feature. Other strong emission lines are N V at 124.0 nm, O IV + Si IV at 140.0 nm, N IV at 148.5 nm, C IV at 155.0 nm, O III at 166.5 nm, N III at 175.0 nm, Si III at 189.2 nm, and C III at 190.8 nm. The observed behavior is quite different from classical symbiotic outbursts, and close monitoring seems in order.”

1993 KA_{2}

Corrigendum. On IAUC 5817, the observer on both dates was T. Gehrels, not D. Rabinowitz.

1993 June 21 Daniel W. E. Green